# Beyond the Surface: Enhancing LLM-as-a-Judge Alignment with Human via Internal Representations

### v1

@@ 5.3 From Tiny to Huge: LAGER Brings Improvements Across Model Scales
The Figure 5 and Figure 7 show the alignment effectiveness of Vanilla Score, E-Score, and lager (w. tuning) across Qwen2.5 models from 0.5B to 72B. Both direct and reasoning conditions show a mostly upward trend in Spearman correlation coefficients as model size increases across Flask, Helpsteer, and BIGGen datasets, with lager leading (e.g., 0.658 in Flask for direct, 0.598 for reasoning). When the model size exceeds 14B, lager achieves greater performance gains over the backbone model than the VScore computed using GPT-4o-mini. Notably, on the BIGGen dataset, lager shows a significant improvement on the 0.5B model scale, elevating the Spearman correlation from a negative value (-0.4 for the Vanilla Score) to a positive 0.1, marking a substantial gain compared to the baseline. These results demonstrate its effectiveness across different model scales and datasets.
@@ 6.1 Instruction Data Selection
Instruction data selection (Cao et al., 2024; Shen, 2024) involves choosing a subset of high-quality data from a given instruction dataset to enhance both the efficiency and performance of the tuning process. In this experiment, we use lager as a scoring metric to select a 10% subset from the alpaca-cleaned-52k dataset [^4] https://huggingface.co/datasets/yahma/alpaca-cleaned (Taori et al., 2023). We compare our subset, Highest lager, with four baselines on finetuning the LLaMA3-8B-base model: Longest instructions, Longest responses (Shen, 2024), Highest VScore, and SuperFiltering (Li et al., 2024d). SuperFiltering is an explicitly designed instruction data filtering method. For methods that require a backbone LLM, we utilize LLaMA3.1-8B-Instruct for a fair comparison. More details are in Appendix F.
@@ B.2 LLM-as-a-Judge
Now there are three different categories of LLM judges: The Point-wise judge model (Kim et al., 2024a; Lin & Chen, 2023; Dong et al., 2024) evaluates individual responses independently, scoring them based on predefined criteria; The pair-wise judge model(Li et al., 2024b; Kim et al., 2024b; Jung et al., 2024) compares two responses directly, determining which one is superior based on a set of evaluation dimensions; The List-wise judge model(Liu et al., 2025) ranks multiple responses in terms of quality, assigning a rank to each based on the overall evaluation. Researchers explore effective approaches to apply LLMs as judges by prompt-based or finetuning-based methods in various scenarios. The prompt-based approaches aim to enhance judgment via step-by-step instructions as well as multi-turn optimizations. G-Eval (Liu et al., 2023) evaluates NLG output quality using Chain-of-Thought (CoT) and Form-Filling paradigms. Portia (Li et al., 2024e) proposes to evaluate the responses with step-by-step reasoning guidelines. Active-critic (Xu et al., 2024) first generates evaluation criteria from data and iteratively refines them. However, the performance of prompt-based LLM evaluators is unsatisfactory (Gu et al., 2024). The finetuning-based approaches further optimize the LLMs with the specialized dataset to adapt the LLMs to judgment tasks either with instruction tuning (Ye & Ng, 2024; Li et al., 2024b) or preference tuning (Ye et al., 2024b; Hu et al., 2024b). InstructScore (Xu et al., 2023) combines human guidance with implicit knowledge from models like GPT-4, finetuned on the LLaMA model, to provide both scores and human-readable diagnostic reports. TIGERScore (Jiang et al., 2023b) is finetuned on the LLaMA-2 model to generate detailed error analysis, helping users understand each identified error and its associated penalty score. Prometheus (Kim et al., 2024a) and Prometheus 2 (Kim et al., 2024b) are finetuned for fine-grained evaluation, achieving new state-of-the-art performance on judgment benchmarks. In this work, we focus on the single-LLM point-wise evaluator without a reference response. Our proposed lager method improves judgment performance by estimation with aggregated layer-wise logits, which is different from the idea of prompt-based and finetuning-based methods.
@@ Appendix C Addition Experimental Results
In this section, Figure 6 shows the comparison of the performance of different scoring methods across models on direct and reasoning conditions. Figure 7 shows the comparison of spearman correlation coefficients for Qwen2.5 models across FLASK, HelpSteer and BIGGen Bench Using Vanilla Score, E-Score, and lager (w. tuning) under the reasoning condition. We present the Spearman correlation coefficients of lager and other baselines for LLaMA3.1-8B-Instruct, Mistral-7B-Instruct-v0.3, and Mistral-Small-24B-Instruct on Flask, HelpSteer, and BiGGen Bench, as shown in Table 5. In Table 6, we present the Pearson correlation coefficients for all models, including lager and other baselines, on the Flask, HelpSteer, and BiGGen Bench.
@@
Table 5: Spearman Correlation Results (Part 2 of 2): Continuation of Table 1. These results are statistically significant, with p-values less than 1e-5. See Table 6 for the full Pearson correlation results.
Model
Flask
Direct
Flask
Reasoning
HelpSteer
direct
HelpSteer
Reasoning
BIGGen Bench
Direct
BIGGen Bench
Reasoning
Average
Fine-tuned Models
TIGERscore-7B - 0.175 - 0.118 - 0.171 0.155
Prometheus2-7B - 0.413 - 0.514 - 0.389 0.448
Close-source Model via API
GPT-4o-mini
Vscore 0.526 0.535 0.482 0.535 0.534 0.509 0.520
E-Score 0.579 0.561 0.500 0.541 0.573 0.530 0.547
Open-source Models
LLaMA3.1-8B-Instruct
GPTScore 0.061 - -0.022 - -0.162 - -0.041
Vscore 0.334 0.429 0.374 0.518 0.273 0.390 0.386
E-Score 0.386 0.446 0.464 0.525 0.352 0.403 0.429
lager (w.o tuning) 0.472 0.456 0.520 0.524 0.475 0.443 0.482
lager (w. tuning) 0.477 0.460 0.515 0.524 0.482 0.444 0.484
Mistral-7B-Instruct-v0.3
GPTScore 0.258 - 0.209 - 0.183 - 0.217
Vscore 0.266 0.269 0.267 0.364 0.138 0.280 0.264
E-Score 0.239 0.279 0.296 0.380 0.185 0.283 0.277
lager (w.o tuning) 0.338 0.295 0.401 0.377 0.353 0.329 0.349
lager (w. tuning) 0.347 0.298 0.403 0.376 0.357 0.333 0.352
LLAMA-3.3-70B-Instruct
GPTScore 0.042 - 0.014 - -0.193 - -0.046
Vscore 0.518 0.567 0.435 0.494 0.559 0.539 0.519
E-Score 0.464 0.540 0.444 0.488 0.530 0.506 0.495
lager (w.o tuning) 0.610 0.598 0.506 0.520 0.597 0.585 0.569
lager (w. tuning) 0.611 0.598 0.504 0.519 0.602 0.584 0.570
@@ D.2 Enhancing LLM in Knowing When They Don’t Know
Knowing when they don’t know, also known as self-knowledge (Yin et al., 2023), is crucial for enhancing the reliability and trustworthiness of LLMs. In this section, we explore whether lager can improve an LLM’s self-knowledge. We evaluate the SelfAware dataset (Yin et al., 2023) using LLaMA3.1-8B-Instruct and Mistral-7B-Instruct-v0.3, which consists of 2337 answerable and 1032 unanswerable questions from five diverse categories. Specifically, given a subjective question in SelfAware, we ask an LLM to evaluate the answerability of this question on a scale of 1 to 5. A score of 1 indicates that the question is completely unanswerable based on the available knowledge, while a score of 5 indicates a highly accurate answer is possible. Following the original paper, we normalize the certainty score to a range of 0-1 and use a threshold of 0.75 as the boundary to classify LLM as either “know” ($\geq$ 0.75) or “don’t know” (<< 0.75).
@@ Appendix F Details of SFT
For lager, E-Score and VScore, we establish seven dimensions (including answer accuracy, logical consistency, relevance, fluency and clarity, length appropriateness, diversity, and instruction difficulty) and utilize the average score of each dimension, evaluated using the LLaMA3.1-8B-Instruct model, for data selection. The specific definitions and scoring standards for these seven dimensions can be found in Table 12. SuperFiltering is an explicitly designed instruction data filtering method. We strictly follow the original paper’s methodology, except for using LLaMA3.1-8B-Instruct as the backbone model for fair comparison. During the SFT training of the Llama-3-8B model, we use 4 L40 GPUs, and the total batch size during the training phase is 64. The specific training parameters and settings are detailed in Table  9.

### v3

@@ 5.3 From Tiny to Huge: LAGER Brings Improvements Across Model Scales
The Figure 5 and Figure 7 show the alignment effectiveness of Vanilla Score, E-Score, and lager (w. tuning) across Qwen2.5 models from 0.5B to 72B. Both direct and reasoning conditions show a mostly upward trend in Spearman correlation coefficients as model size increases across Flask, Helpsteer, and BIGGen datasets, with lager leading (e.g., 0.658 in Flask for direct, 0.598 for reasoning). When the model size exceeds 14B, lager achieves greater performance gains over the backbone model than the VScore computed using GPT-4o-mini. Notably, on the BIGGen dataset, lager shows a significant improvement on the 0.5B model scale, elevating the Spearman correlation from a negative value (-0.4 for the Vanilla Score) to a positive 0.1, marking a substantial gain compared to the baseline. These results demonstrate its effectiveness across different model scales and datasets, and the benefits brought by lager are scale-transferable and synergistically amplified as model capacity grows.
@@ 6.1 Instruction Data Selection
Instruction data selection (Cao et al., 2024; Shen, 2024) involves choosing a subset of high-quality data from a given instruction dataset to enhance both the efficiency and performance of the tuning process. In this experiment, we use lager as a scoring metric to select a 10% subset from the alpaca-cleaned-52k dataset [^4] https://huggingface.co/datasets/yahma/alpaca-cleaned (Taori et al., 2023). We compare our subset, Highest lager, with four baselines on finetuning the LLaMA3-8B-base model: Longest instructions, Longest responses (Shen, 2024), Highest VScore, and SuperFiltering (Li et al., 2024e). SuperFiltering is an explicitly designed instruction data filtering method. For methods that require a backbone LLM, we utilize LLaMA3.1-8B-Instruct for a fair comparison. More details are in Appendix F.
@@ 7 Conclusion
Acknowledgements
This project was supported by National Natural Science Foundation of China (No. 62306132), Guangdong Basic and Applied Basic Research Foundation (No. 2025A1515011564), Natural Science Foundation of Shanghai (No. 25ZR1402136). We thank the anonymous reviewers for their insightful feedback on this work.
@@ B.2 LLM-as-a-Judge
Now there are three different categories of LLM judges: The Point-wise judge model (Kim et al., 2024a; Lin & Chen, 2023; Dong et al., 2024) evaluates individual responses independently, scoring them based on predefined criteria; The pair-wise judge model(Li et al., 2024b; Kim et al., 2024b; Jung et al., 2024) compares two responses directly, determining which one is superior based on a set of evaluation dimensions; The List-wise judge model(Liu et al., 2025) ranks multiple responses in terms of quality, assigning a rank to each based on the overall evaluation. Researchers explore effective approaches to apply LLMs as judges by prompt-based or finetuning-based methods in various scenarios. The prompt-based approaches aim to enhance judgment via step-by-step instructions as well as multi-turn optimizations. G-Eval (Liu et al., 2023) evaluates NLG output quality using Chain-of-Thought (CoT) and Form-Filling paradigms. Portia (Li et al., 2024f) proposes to evaluate the responses with step-by-step reasoning guidelines. Active-critic (Xu et al., 2024) first generates evaluation criteria from data and iteratively refines them. However, the performance of prompt-based LLM evaluators is unsatisfactory (Gu et al., 2024). The finetuning-based approaches further optimize the LLMs with the specialized dataset to adapt the LLMs to judgment tasks either with instruction tuning (Ye & Ng, 2024; Li et al., 2024b) or preference tuning (Ye et al., 2024b; Hu et al., 2024b). InstructScore (Xu et al., 2023) combines human guidance with implicit knowledge from models like GPT-4, finetuned on the LLaMA model, to provide both scores and human-readable diagnostic reports. TIGERScore (Jiang et al., 2023b) is finetuned on the LLaMA-2 model to generate detailed error analysis, helping users understand each identified error and its associated penalty score. Prometheus (Kim et al., 2024a) and Prometheus 2 (Kim et al., 2024b) are finetuned for fine-grained evaluation, achieving new state-of-the-art performance on judgment benchmarks. In this work, we focus on the single-LLM point-wise evaluator without a reference response. Our proposed lager method improves judgment performance by estimation with aggregated layer-wise logits, which is different from the idea of prompt-based and finetuning-based methods.
@@ Appendix C Addition Experimental Results
In this section, Figure 6 shows the comparison of the performance of different scoring methods across models on direct and reasoning conditions. Figure 7 shows the comparison of spearman correlation coefficients for Qwen2.5 models across FLASK, HelpSteer and BIGGen Bench Using Vanilla Score, E-Score, and lager (w. tuning) under the reasoning condition. In Table 5, we present the Pearson correlation coefficients for all models, including lager and other baselines, on the Flask, HelpSteer, and BiGGen Bench.
@@ D.2 Enhancing LLM in Knowing When They Don’t Know
Knowing when they don’t know, also known as self-knowledge (Yin et al., 2023), is crucial for enhancing the reliability and trustworthiness of LLMs. In this section, we explore whether lager can improve an LLM’s self-knowledge. We evaluate the SelfAware dataset (Yin et al., 2023) using LLaMA3.1-8B-Instruct and Mistral-7B-Instruct-v0.3, which consists of 2337 answerable and 1032 unanswerable questions from five diverse categories. Specifically, given a subjective question in SelfAware, we ask an LLM to evaluate the answerability of this question on a scale of 1 to 5. A score of 1 indicates that the question is completely unanswerable based on the available knowledge, while a score of 5 indicates a highly accurate answer is possible. Following the original paper, we normalize the certainty score to a range of 0-1 and use a threshold of 0.75 as the boundary to classify LLM as either "know" ($\geq$ 0.75) or "don’t know" (<< 0.75).
@@ Appendix F Details of SFT
For lager, E-Score and VScore, we establish seven dimensions (including answer accuracy, logical consistency, relevance, fluency and clarity, length appropriateness, diversity, and instruction difficulty) and utilize the average score of each dimension, evaluated using the LLaMA3.1-8B-Instruct model, for data selection. The specific definitions and scoring standards for these seven dimensions can be found in Table 12. SuperFiltering is an explicitly designed instruction data filtering method. We strictly follow the original paper’s methodology, except for using LLaMA3.1-8B-Instruct as the backbone model for fair comparison. During the SFT training of the Llama-3-8B model, we use 4 L40 GPUs, and the total batch size during the training phase is 64. The specific training parameters and settings are detailed in Table  8.